A student throws a water balloon vertically downward from the top of the building. The balloon leaves the thrower’s hand with a speed of 6.0m/s. Air resistance may be ignored so the water balloon is in free fall after it leaves the thrower’s hand.


A.) What is it’s speed after falling 2.0s?


B.) How far does it fall in 2s?


C.) What is the magnitude of it’s velocity after falling 10m?




1
Expert's answer
2021-03-26T11:42:04-0400

Answer

Using newton equation of motion

A. Velocity after t time

v=u+gt

"v=0+9.8\times2=19.6m/s"

B. Distance travelled after t=2sec

"S=ut+0.5gt^2\\=0+0.5\times9.8\times4\\=19.6m"

C. Velocity after t=10sec


"v=0+9.8\times10=98m/s"